OpenStudy (anonymous):

f(x)=(2x+3)(x-2) find the vertex

OpenStudy (anonymous):

I got (1/4,-95/16) but that's not right .

OpenStudy (anonymous):

multiple choice?

OpenStudy (anonymous):

I have to graph it

OpenStudy (amoodarya):

turn it to standard form y=a(x-)^2+c so x=b is x of vertex put it ant find y of vertex

OpenStudy (amoodarya):

\[f(x)=(2x+3)(x-2) = 2(x+\frac{ 3 }{ 2})(x-2)=2(x^2-\frac{ 1 }{ 2}x-3)\] \[2(x^2-\frac{ 1 }{ 2}x-3)=2( (x-\frac{ 1}{4})^2-(\frac{ 1}{4})^2-3)=2(x-\frac{ 1}{4})^2-2((\frac{ 1}{4})^2-3)\]

OpenStudy (amoodarya):

so x of vertex is x=1/4